Help needed on these questions. I'm having trouble understanding the concept and a clear explanation would be very helpful. Thank you very much!

In Japan, one worker can make 5 tons of rubber or 80 radios. In Malaysia, one worker can make 10 tons of rubber or 40 radios. Each country has 10 workers.

Who has the absolute advantage in producing radios?

Who has the comparative advantage in producing rubber?

What should Japan specialize in producing to benefit from trade?

What is the relationship between quantity demanded and quantity supplied at equilibrium?

What is the relationship when there is a shortage?

What is the relationship when there is a surplus?
Answer #1
Japan Malaysia
Rubber 5 10
Radio 80 40

1. Japan has an absolute advantage in producing radios because Japan can produce more radio by using same labor than Malaysia.

2. Japan's opportunity cost for producing rubber = 80/5 = 16 radios

And Malaysia's opportunity cost for producing rubber= 40/10= 4 radios.

Because Malaysia has a lower opportunity cost for producing rubber. This implies that Malaysia has a comparative advantage in producing rubber and completely specialize in producing rubber.

3. And Japan has a comparative advantage in producing radios, so Japan should specialize in producing radios to benefit from trade.

4. At equilibrium quantity demanded equals quantity supplied. The quantity demanded at a particular price exactly met the quantity supplied , which implies equilibrium.

5. When there is a shortage , quantity demanded exceeds quantity supplied.

6. When there is a surplus . quantity supplied exceeds quantity demanded.